Please help... GRE cramming and struggling.
7x+3y=12
3x+7y=6
x-y=?
Have tried several methods, keep getting crazy fractions.
Thanks so much.
E

Answer by mananth(16949)

7 x + 3 y = 12 .............1
3 x + 7 y = 6 .............2
Eliminate y
multiply (1)by 7
Multiply (2) by -3
49 x 21 y = 84
-9 x -21 y = -18
Add the two equations
40 x = 66
/ 40
x = 1.65
plug value of x in (1)
7 x + 3 y = 12
11.55 + 3 y = 12
3 y = 12 -11.55
3 y = 0.45
y = 0.15
x-y = 1.5
m.ananth@hotmail.ca